OpenStudy (anonymous):

Determine whether the graph of y = x2 + 2x − 8 has a maximum or minimum point, then find the maximum or minimum value.

OpenStudy (cwrw238):

as the coefficient of x is positive it has a minimum do you know calculus?

OpenStudy (anonymous):

no

OpenStudy (cwrw238):

oh ok we need to convert to the form (x - a)^2 + b where the mimimum value will be b

OpenStudy (cwrw238):

y = x^2 + 2x − 8 = ( x + 1)^ 2 - 1 - 8 = (x + 1)^2 - 9 giving a minimum value of -9

OpenStudy (cwrw238):

http://www.wolframalpha.com/input/?i=x%5E2+%2B+2x+-+8+%3D+0 shows the graph of the function

OpenStudy (cwrw238):

have you any more questions?

OpenStudy (anonymous):

Maximum; (-1, -9) Minimum; (-1, -9) Maximum; (-9, -1) Minimum; (-9, -1)

OpenStudy (cwrw238):

what you have there is the coordinates of the minimum value which occurs when x+ 1 = 0 or x = -1 the correct option is the second one